A horizontal force of 930 N is needed to overcome
Answer #1

Considering forces in vertical direction, weight of the crate = normal force on crate by floor

Mg = N

Consider forces in horizontal direction, applied force = force of friction

Fapp= fs

Fapp= 930N

Force of static friction ,fs = ms N , where ms is the coefficient of static friction

F= ms N

F = ms Mg

ms = F/ Mg = 930 / ( 150 x 9.8) = 0.633